Intermediate Microeconomics

2. The following are the demand and supply functions of a hypothetical product.

Qd = -.6P + 8

Qs = .4P -2

(a) Find the equilibrium price and quantity.

(b) Calculate the price elasticity of

(i) demand and

(ii) supply at the equilibrium point.
Answer #1

Qued we have, - 0.6 P + 8 @ P. Ovup-2 for Equilibrium in the porket od=8 -06 P + 8 = oup-2 878 - 04 +0.6P 10 - 10 p 100 cp 10Therefore, Elasticity of sermond am fx ed = fx (0-6) . ed at Equilibrium Poice & dontity - ed = 10 x test Ted =-3 / Elasticit